The number of accidents in a highway as recorded every month over a 9 month period are 15,18,9,11,14,10,8,13,19. Test at 5% frequencies are in in agreement with the belief that the number of accidents was the same during the 9 months. it is given that the table values of x² at 5% level for 8 d.o.f. are 15.5 and 16.9 respectively.


1
Expert's answer
2021-05-07T10:37:13-0400

Let,

"H_o:" The number of accidents ocur during 9 months are same with expected frequencies "(E_i=\dfrac{N}{9}=\dfrac{117}{9}=13)"


The distribution table is-






The value of "\chi^2=\sum{\dfrac{(O_i-E_i)^2}{E_i}}=\dfrac{100}{13}=7.6932"


The tabulated value of "\chi^2" at 8 d.o.f. is 16.9


Conclusion: As the calculated value of "\chi^2" is less than the standard value so, Null hypothesis is accepted i.e. Number of accidents are same.